On 1.11.2019 you enter into a long futures contract for 100 ounces of gold at $1,500 an ounce. Your initial margin is set at 15% of the contract value. The settlement prices for gold on the subsequent two days are: $1550 on 2.11.2019 and $1,470 on 3.11.2019. Show the “marking-to-market” and the adjustment in the margin account at the end of each of the two trading days. What is the total 2-day gain or loss on the position?
Answer #1

Contract size = 100 ounce of gold      
Position = long      
Buying or long price =    1500  
Closing or sale price on day 1=   1550  
      
So profit or loss on day 1 = (1550-1500)*100      
5000      
      
Initial margin = 15% of (100*1500)=   22500  


So balance margin on day 1= 22500+5000=      
$22,500.00      


So Adjustment not required.       
So MTM on day 1 is $5000. Adjustment to margin is 0      


profit or loss on day 2 = (1470-1550)*100      
-8000      
      
margin Balance=    $22,500.00  
So balance margin on day 2= 27500-8000=   $14,500.00  
      
Balance is below $22500 initila margin. so Adjustment to margin account or margin call will made= 22500-19500=   $8,000.00  
      
MTM on day 2 is -$8000. Adjustment required is $3000